OpenStudy (anonymous):

The measure of an angle is 40 degrees less than four times the measure of its complement. Find the measure of the angle.

OpenStudy (anonymous):

Call x the angle. Complement of x: 4(90 - x) x = 4(90 - x) - 40 = -4x + 320 x = ...

OpenStudy (anonymous):

can you solve it all the way

OpenStudy (anonymous):

@thu1935

OpenStudy (whpalmer4):

Actually, if \(x\) is the angle, the complement is \(90-x\), not \(4(90-x)\) Subsequent lines appear correct, however.

OpenStudy (whpalmer4):

\[x = 4(90-x) -40\]@cutegirl What do you get if you use the distributive property to remove the parentheses?